# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ 5 Conclusion
This research represents a thorough benchmark of machine and deep learning methodologies for tree species classification, leveraging multispectral ALS data. By employing both high-density HeliALS data ($>1000$ $\mathrm{pts}/\mathrm{m}^{2}$) and lower-density Optech Titan data (35 $\mathrm{pts}/\mathrm{m}^{2}$), we have provided a comprehensive comparison of various algorithmic approaches. For the benchmark study, we established a training set of 1065 tree segments and a test set of 5261 segments representing nine species in Southern Finland. In our benchmark, point-based deep learning methods outperformed both image-based deep learning and machine learning methods on the high-density HeliALS data, while the best point-based deep learning methods performed equivalently with machine learning methods on the lower-density Optech Titan data. On the HeliALS data, a Point Transformer model topped the list with an overall (macro-average) accuracy of 87.9% (74.5%). On the Optech Titan dataset, a random forest classifier reached the highest overall (macro-average) accuracy of 79.9% (57.6%).
@@
Furthermore, we show that the accuracy improvement provided by deep learning methods increases with training set size as discussed in Sec. 4.2.1. We observe that the classification error obeys a power law $\varepsilon(m)=Am^{-\alpha}$ with respect to the training set size both for a point transformer model and a random forest classifier. Importantly, the scaling exponent $\alpha$ is approximately twice higher for the deep learning model leading to a drastically faster convergence of the classification error. Based on our analysis, the studied deep learning method outperforms the random forest classifier for training set sizes exceeding a few hundred trees on both datasets. Using the fitted scaling laws on the HeliALS data, we predict that a macro-average accuracy of 90% would require 14 000 training trees with the point transformer model and $4.9\times 10^{6}$ training trees with the random forest classifier, thus highlighting the benefit of deep learning approaches on large datasets. The development of the crowdsourced field reference dataset, along with the innovative browser-based crowdsourcing tool, represent a significant contribution to the scientific community aiding the collection of large datasets required to scale up these methods to larger forest areas. These resources will facilitate future research and development in this domain.
The ability to accurately identify tree species at the individual tree level is essential for sustainable multifunctional forest management practices. Future research should focus on further refining deep learning architectures for multispectral ALS data, and addressing the challenges associated with identifying rare species. More detailed studies are also needed on the effects of different preprocessing steps and data augmentation methods on the performance of deep learning methods. Integration of tree species classification to direct measurements of stem curve and volume using a high-resolution airborne laser scanning system, as depicted in (Hyyppä et al., 2022), would be further step in developing a fully automatic reference collection technique for forest inventory. Additionally, investigating the transferability of these models across different geographic regions and sensor platforms is crucial for broader adoption. From the practical point of view, it is also important to pay attention to datasets that are available nationwide or over large areas.
@@ Acknowledgements
We gratefully acknowledge the Research Council of Finland (RCF), CHIST-ERA, the NextGenerationEU instrument and Finnish Government (VN) funding for the following projects “Mapping of forest health, species and forest fire risks using Novel ICT Data and Approaches” (RCF 344755, CHIST-ERA-19-CES-001, main project for this study until March 2024), “Capturing structural and functional diversity of trees and tree communities for supporting sustainable use of forests” (RCF 348644, NextGenerationEU), “Forest-Human-Machine Interplay – Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences” (RCF 359175), “Collecting accurate individual tree information for harvester operation decision making” (RCF 359554), “High-performance computing allowing high-accuracy country-level individual tree carbon sink and biodiversity mapping” (RCF 359203), “Measuring Spatiotemporal Changes in Forest Ecosystem” (decision number 346382)”, “Digital technologies, risk management solutions and tools for mitigating forest disturbances” (RCF 353264, NextGenerationEU), and the Ministry of Agriculture and Forestry grant “Future Forest Information System at Individual Tree Level” (VN/3482/2021). M. Myllymäki was funded by the European Union – NextGenerationEU in RCF project "Artificial intelligence, spatial statistics and Earth observation for digital twinning of forest diversity" (RCF 348154) under the Research Council of Finland’s flagship ecosystem for Forest–Human–Machine Interplay—Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences (UNITE) (357909). J. Kostensalo was funded by "Digital Twin of Boreal Forest Structure with Remote Sensing and Forest Inventories" (RF 361209). S. Puliti and M. Astrup conducted the work as part of the Center for Research-based Innovation SmartForest: Bringing Industry 4.0 to the Norwegian forest sector (NFR SFI project no. 309671, smartforest.no).
@@
Appendix A. Macro-average accuracy by profile category and tree crown category
@@
Figure 15: Macro-average accuracy by profile category for the best machine learning and deep learning methods for HeliALS (Panel A) and Optech Titan (Panel B) datasets. Median values of macro-average accuracy were calculated across the methods that utilize the same dataset and approach type (DL/ML). The best method was determined by macro $F_{1}$ score. (a) For the HeliALS dataset, the best deep learning method is FGI-PointTransformerWeighted-DL-3D (Macro $F_{1}=0.73$) and the best machine learning method is FGI-RF-ML (Macro $F_{1}=0.64$). (b) For the Optech Titan dataset, the best deep learning method is FGI-PointTransformer-DL-3D (Macro $F_{1}=0.58$) and the best machine learning method is IBL-BalancedRF-ML (Macro $F_{1}=0.57$).
@@
Figure 16: Macro-average accuracy by crown class category for the best machine learning and deep learning methods for HeliALS (Panel A) and Optech Titan (Panel B) datasets. Median values of macro-average accuracy were calculated across the methods that utilize the same dataset and approach type (DL/ML). The best method was determined by macro $F_{1}$ score. (a) For the HeliALS dataset, the best deep learning method is FGI-PointTransformerWeighted-DL-3D (Macro $F_{1}=0.73$) and the best machine learning method is FGI-RF-ML (Macro $F_{1}=0.64$). (b) For the Optech Titan dataset, the best deep learning method is FGI-PointTransformer-DL-3D (Macro $F_{1}=0.58$) and the best machine learning method is IBL-BalancedRF-ML (Macro $F_{1}=0.57$).
@@
Figures 15 and 16 present the macro-average accuracies by profile category and by crown class category, respectively, for the best methods according to the macro $F_{1}$ score. The figures have been made for comparison against Fig. 12 and Fig. 13, which present the overall accuracies for the same categories. Notice the different scaling of the y-axis compared to the Fig. 12 and Fig. 13.
Appendix B. Confusion matrices
@@
Tables 613 show the normalized confusion matrices of the deep learning methods that use the HeliALS dataset, with the bolded diagonal entries representing the per-class recall. Furthermore, Tables 1418 show the normalized confusion matrices of the deep learning methods that use the Optech Titan dataset, and Tables 1926 show the normalized confusion matrices of the machine learning methods that use the Optech Titan dataset.
Table 6: Confusion matrix: FBK-PointNet++-DL-3D, HeliALS
@@
Table 7: Confusion matrix: FGI-Point2Vec-DL-3D, HeliALS
@@
Table 8: Confusion matrix: FGI-DGCNN-DL-3D, HeliALS
@@
Table 9: Confusion matrix: FGI-PointNet-DL-3D, HeliALS
@@
Table 10: Confusion matrix: FGI-PointTransformer-DL-3D, HeliALS
@@
Table 11: Confusion matrix: SA-Detailview-DL-2D, HeliALS
@@
Table 12: Confusion matrix: SLU-YOLOv8-DL-2D, HeliALS
@@
Table 13: Confusion matrix: TUW-PointNet++-DL-3D, HeliALS
@@
Table 14: Confusion matrix: FBK-PointNet++-DL-3D, Optech Titan
@@
Table 15: Confusion matrix: FGI-DGCNN-DL-3D, Optech Titan
@@
Table 16: Confusion matrix: FGI-PointNet-DL-3D, Optech Titan
@@
Table 17: Confusion matrix: FGI-PointTransformerWeighted- DL-3D, Optech Titan
@@
Table 18: Confusion matrix: NTNU-ConvNeXt-T-DL-2D, Optech Titan
@@
Table 19: Confusion matrix: Aalto-RF-ML, Optech Titan
@@
Table 20: Confusion matrix: FGI-RF-ML, Optech Titan
@@
Table 21: Confusion matrix: LUKE-MultiRF-ML, Optech Titan
@@
Table 22: Confusion matrix: IDEAS-RF-ML, Optech Titan
@@
Table 23: Confusion matrix: UEF-LGBM-ML, Optech Titan
@@
Table 24: Confusion matrix: UEF-RF-ML, Optech Titan
@@
Table 25: Confusion matrix: UEF-SVM-ML, Optech Titan
@@
Table 26: Confusion matrix: UPV-GB-ML, Optech Titan

### v2

@@ 5 Conclusion
This research represents a thorough benchmark of machine and deep learning methodologies for tree species classification, leveraging multispectral ALS data. By employing both high-density HeliALS data (>1000>1000 $\mathrm{pts}/\mathrm{m}^{2}$) and lower-density Optech Titan data (35 $\mathrm{pts}/\mathrm{m}^{2}$), we have provided a comprehensive comparison of various algorithmic approaches. For the benchmark study, we established a training set of 1065 tree segments and a test set of 5261 segments representing nine species in southern Finland. In our benchmark, point-based deep learning methods outperformed both image-based deep learning and machine learning methods on the high-density HeliALS data, while the best point-based deep learning methods performed equivalently with machine learning methods on the lower-density Optech Titan data. On the HeliALS data, a Point Transformer model topped the list with an overall (macro-average) accuracy of 87.9% (74.5%). On the Optech Titan dataset, a random forest classifier reached the highest overall (macro-average) accuracy of 79.9% (57.6%), though three other methods were within error bars in terms of the overall accuracy.
@@
Furthermore, we show that the accuracy improvement provided by deep learning methods increases with training set size as discussed in Sec. 4.4.1. We observe that the classification error obeys a power law $\varepsilon(m)=Am^{-\alpha}$ with respect to the training set size both for a point transformer model and for a random forest classifier. Importantly, the scaling exponent $\alpha$ is approximately twice higher for the deep learning model leading to a drastically faster convergence of the classification error. Based on our analysis, the studied deep learning method outperforms the random forest classifier for training set sizes exceeding a few hundred segments on both datasets. Using the fitted scaling laws on the HeliALS data, we predict that a macro-average accuracy of 90% would require 14 000 training segments with the point transformer model and $4.9\times 10^{6}$ training segments with the random forest classifier, thus highlighting the benefit of deep learning approaches on large datasets. The development of the crowdsourced field reference dataset, along with the innovative browser-based crowdsourcing tool, represent a significant contribution to the scientific community aiding the collection of large datasets required to scale up these methods to larger forest areas. These resources will facilitate future research and development in this domain.
The ability to accurately identify tree species at the individual tree level is essential for sustainable municipality and multifunctional forest management practices. Future research should focus on further refining deep learning architectures for multispectral ALS data, and addressing the challenges associated with identifying rare species. More detailed studies are also needed on the effects of different preprocessing steps and data augmentation methods on the performance of deep learning methods. Integration of tree species classification to direct measurements of stem curve and volume using a high-resolution airborne laser scanning system, as depicted in (hyyppa2022direct), would be further step in developing a fully automatic reference collection technique. Additionally, investigating the transferability of these models across different geographic regions and sensor platforms is crucial for broader adoption. From the practical point of view, it is also important to pay attention to datasets that are available nationwide or over large areas.
Data availability
We have made available the Espoonlahti dataset for tree species classification in Zenodo: https://doi.org/10.5281/zenodo.17077255. The dataset includes both HeliALS and Optech Titan point clouds for the studied tree segments together with ground-truth species labels. The published Zenodo dataset will be described in more detail in a forthcoming data manuscript.
@@ Acknowledgements
We gratefully acknowledge the Research Council of Finland (RCF), CHIST-ERA, the NextGenerationEU instrument and Finnish Government (VN) funding for the following projects “Mapping of forest health, species and forest fire risks using Novel ICT Data and Approaches” (RCF 344755, CHIST-ERA-19-CES-001, main project for this study until March 2024), “Capturing structural and functional diversity of trees and tree communities for supporting sustainable use of forests” (RCF 348644, NextGenerationEU), “Forest-Human-Machine Interplay – Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences” (RCF 359175), “Collecting accurate individual tree information for harvester operation decision making” (RCF 359554), “High-performance computing allowing high-accuracy country-level individual tree carbon sink and biodiversity mapping” (RCF 359203), “Measuring Spatiotemporal Changes in Forest Ecosystem” (decision number 346382)”, “Digital technologies, risk management solutions and tools for mitigating forest disturbances” (RCF 353264, NextGenerationEU), and the Ministry of Agriculture and Forestry grant “Future Forest Information System at Individual Tree Level” (VN/3482/2021). M. Myllymäki was funded by the European Union – NextGenerationEU in RCF project "Artificial intelligence, spatial statistics and Earth observation for digital twinning of forest diversity" (RCF 348154) under the Research Council of Finland’s flagship ecosystem for Forest–Human–Machine Interplay—Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences (UNITE) (357909). J. Kostensalo was funded by "Digital Twin of Boreal Forest Structure with Remote Sensing and Forest Inventories" (RCF 361209). S. Puliti and M. Astrup conducted the work as part of the Center for Research-based Innovation SmartForest: Bringing Industry 4.0 to the Norwegian forest sector (NFR SFI project no. 309671, smartforest.no).
@@
Appendix A Distribution of mean intensity for the studied tree species
Multispectral data provides a benefit for tree species classification since different tree species have a different reflectance response as a function of wavelength (hovi2017spectral). Figures A.1 and A.2 illustrate the species-wise distribution of mean intensity for the three channels of the HeliALS system and the Optech system, respectively. For example, Fig. A.1 shows that the mean intensity of the 3rd channel ($\lambda=532$ nm) of the HeliALS data differs significantly between pine and birch, while the difference of mean intensities is smaller for these two species at the other two wavelengths. Overall, the variations in the intensity distributions between different species showcase the advantage of multispectral ALS for tree species classification.
Figure A.1: Distribution of mean intensity for the three channels of the HeliALS system for the studied nine species across all segments of our dataset. Here, blue corresponds to the VQ-840-G scanner, green to the miniVUX-1DL scanner, and red to the VUX-1HA scanner.
Figure A.2: Distribution of mean intensity for the three channels of the Optech Titan system for the studied nine species across all segments of our dataset. Here, blue corresponds to $\lambda=532$ nm, green to $\lambda=1064$ nm, and red to $\lambda=1550$ nm.
Appendix B Comparison of algorithm running times
The computational cost of the FGI-RF-ML, the FGI-PointTransformer-DL-3D and the DetailView-DL-2D methods was assessed to determine their scalability to larger datasets. The comparison was performed separately for model training phases and inference phases. The training set contained 1065 samples, while the test set comprised 5261 samples. In addition, the running times of the FGI-RF-ML and the FGI-PointTransformer-DL-3D were reported both for the Optech Titan dataset and for the HeliALS dataset. The hardware specifications of the experiments are listed in Table B.1. The FGI-RF-ML method was run on computer 1, and the FGI-PointTransformer-DL-3D and the DetailView-DL-2D methods on computer 2.
Table B.1: Hardware specifications of the two computers used for runtime measurements.
Specification Computer 1 Computer 2
CPU Intel Core i7-9850H (6 cores, 2.6 GHz) Intel Xeon(R) Gold 6244 (16 cores, 3.6 GHz)
RAM 32 GB 512 GB
GPU not utilized NVIDIA RTX A6000 (48 GB)
Operating System Windows 10 Ubuntu 22.04.5 LTS
Figure B.1: Comparison of running times between three tree species classification methods in different datasets. The training set size was 1065 tree segments and the test set size was 5261 tree segments. The dashed horizontal line represents the total running time including feature computation and classification. For deep learning based methods, the feature computation is inherent to the computations and has not been separately reported.
The results of the running time comparison have been visualized in Fig. B.1. In the training phase, the computational cost of the FGI-RF-ML and the FGI-PointTransformer-DL-3D methods are quite similar. However, the DetailView-DL-2D method has a substantionally longer training time than either of the previous methods. The FGI-RF-ML method is the fastest to train and can be trained in less than six minutes on our Optech Titan dataset.
During inference, the FGI-PointTransformer-DL-3D method performs the best in terms of running time. When using the HeliALS dataset, the FGI-RF-ML method is approximately 177 times slower than the FGI-PointTransformer-DL-3D method. Similarly, the inference speed of the DetailView-DL-2D method is approximately 52 times slower when compared to the FGI-PointTransformer-DL-3D method. The FGI-RF-ML method uses a majority of the computation time for the calculation of the point cloud features (99,98% of the total time for the HeliALS dataset and 99,74% of the total time for the Optech Titan dataset), whereas the prediction process of the RF classifier is relatively fast. It can be hypothesized that a more efficient implementation of the feature calculation process could substantially increase the inference speed of FGI-RF-ML.
The implementation details of the inference process influence the results to a certain degree. For example, the FGI-PointTransformer-DL-3D was trained separately for five times and the five model instances were used to infer the tree species on the test dataset using a majority voting scheme. Likewise, the DetailView-DL-2D method performed the inference for 50 randomly augmented versions of each test set sample obtaining the prediction by selecting the species with the maximum cumulative prediction probability. When these factors are taken into consideration, the FGI-PointTransformer-DL-3D method is approximately five times faster than the DetailView-DL-2D method.
@@ Appendix B Comparison of algorithm running times
Appendix C Macro-average accuracy by profile category and tree crown category
@@
Figure C.1: Macro-average accuracy by profile category for the best machine learning and deep learning methods for HeliALS (Panel A) and Optech Titan (Panel B) datasets. Median values of macro-average accuracy were calculated across the methods that utilize the same dataset and approach type (DL/ML). The best method was determined by macro $F_{1}$ score. (a) For the HeliALS dataset, the best deep learning method is FGI-PointTransformerWeighted-DL-3D (Macro $F_{1}=0.73$) and the best machine learning method is FGI-RF-ML (Macro $F_{1}=0.64$). (b) For the Optech Titan dataset, the best deep learning method is FGI-PointTransformer-DL-3D (Macro $F_{1}=0.58$) and the best machine learning method is IBL-BalancedRF-ML (Macro $F_{1}=0.57$).
@@
Figure C.2: Macro-average accuracy by crown class category for the best machine learning and deep learning methods for HeliALS (Panel A) and Optech Titan (Panel B) datasets. Median values of macro-average accuracy were calculated across the methods that utilize the same dataset and approach type (DL/ML). The best method was determined by macro $F_{1}$ score. (a) For the HeliALS dataset, the best deep learning method is FGI-PointTransformerWeighted-DL-3D (Macro $F_{1}=0.73$) and the best machine learning method is FGI-RF-ML (Macro $F_{1}=0.64$). (b) For the Optech Titan dataset, the best deep learning method is FGI-PointTransformer-DL-3D (Macro $F_{1}=0.58$) and the best machine learning method is IBL-BalancedRF-ML (Macro $F_{1}=0.57$).
@@
Figures C.1 and C.2 present the macro-average accuracies by profile category and by crown class category, respectively, for the best methods according to the macro $F_{1}$ score. The figures have been made for comparison against Fig. 12 and Fig. 13, which present the overall accuracies for the same categories. Notice the different scaling of the y-axis compared to the Fig. 12 and Fig. 13.
Appendix D Confusion matrices for all methods and confusion matrices for best methods with raw counts
Figure D.1: (a) Unnormalized confusion matrix of FGI-RF-ML obtaining the highest macro $F_{1}$ score (0.64) among the machine learning models using HeliALS data. FGI-RF-ML produced 55 missing predictions. (b) Unnormalized confusion matrix of FGI-PointTransformerWeighted-DL-3D obtaining the highest macro $F_{1}$ score (0.73) among the deep learning models using HeliALS data. (c) Unnormalized confusion matrix of IBL-BalancedRF-ML obtaining the highest macro $F_{1}$ score (0.57) among the machine learning models using Optech Titan data. (d) Unnormalized confusion matrix of FGI-PointTransformer-DL-3D obtaining the highest macro $F_{1}$ score (0.58) among the deep learning models using Optech Titan data. All confusion matrices show raw counts. Precisions are displayed below the matrix. Row and column sums are shown to the right and bottom of the matrix in gray.
Figure D.1 shows the raw confusion matrices of counts for the best performing machine learning and deep learning methods on the HeliALS and Optech Titan datasets. The corresponding normalized confusion matrices are shown in Fig. 7.
@@ Appendix D Confusion matrices for all methods and confusion matrices for best methods with raw counts
Tables D.1D.8 show the row-normalized confusion matrices of the deep learning methods using the HeliALS dataset, with the bolded diagonal entries representing the per-class recall. Furthermore, Tables D.9D.13 show the normalized confusion matrices of the deep learning methods using the Optech Titan dataset, and Tables D.14D.21 show the normalized confusion matrices of the machine learning methods using the Optech Titan dataset.
Table D.1: Confusion matrix: FBK-PointNet++-DL-3D, HeliALS
@@
Table D.2: Confusion matrix: FGI-Point2Vec-DL-3D, HeliALS
@@
Table D.3: Confusion matrix: FGI-DGCNN-DL-3D, HeliALS
@@
Table D.4: Confusion matrix: FGI-PointNet-DL-3D, HeliALS
@@
Table D.5: Confusion matrix: FGI-PointTransformer-DL-3D, HeliALS
@@
Table D.6: Confusion matrix: SA-Detailview-DL-2D, HeliALS
@@
Table D.7: Confusion matrix: SLU-YOLOv8-DL-2D, HeliALS
@@
Table D.8: Confusion matrix: TUW-PointNet++-DL-3D, HeliALS
@@
Table D.9: Confusion matrix: FBK-PointNet++-DL-3D, Optech Titan
@@
Table D.10: Confusion matrix: FGI-DGCNN-DL-3D, Optech Titan
@@
Table D.11: Confusion matrix: FGI-PointNet-DL-3D, Optech Titan
@@
Table D.12: Confusion matrix: FGI-PointTransformerWeighted- DL-3D, Optech Titan
@@
Table D.13: Confusion matrix: NTNU-ConvNeXt-T-DL-2D, Optech Titan
@@
Table D.14: Confusion matrix: Aalto-RF-ML, Optech Titan
@@
Table D.15: Confusion matrix: FGI-RF-ML, Optech Titan
@@
Table D.16: Confusion matrix: LUKE-MultiRF-ML, Optech Titan
@@
Table D.17: Confusion matrix: IDEAS-RF-ML, Optech Titan
@@
Table D.18: Confusion matrix: UEF-LGBM-ML, Optech Titan
@@
Table D.19: Confusion matrix: UEF-RF-ML, Optech Titan
@@
Table D.20: Confusion matrix: UEF-SVM-ML, Optech Titan
@@
Table D.21: Confusion matrix: UPV-GB-ML, Optech Titan